Bob runs 25m straight on a road in 3.5 sec. He is then hit and pushed 4.2 m backwards in 2.3 sec. He again runs straight forward another 32m in 7.5 sec. Calculate his average velocity for (a) each of the 3 intervals and (b) for the entire motion
1
Expert's answer
2020-05-04T12:49:33-0400

a)

"v_1=\frac{25}{3.5}=7.1\frac{m}{s}"

"v_2=\frac{-4.2}{2.3}=-1.8\frac{m}{s}"

"v_3=\frac{32}{7.5}=4.3\frac{m}{s}"

b)


"v=\frac{25-4.2+32}{3.5+2.3+7.5}=4.0\frac{m}{s}"